Use a graphing utility to find or approximate solution(s) of the given equation.
20x+5=24x^2+6x

Answer by addingup(3677)

Solve for x:
20x+5= 24x^2+6x
Subtract 24x^2+6x from both sides:
-24x^2+14x+5= 0
Factor the left side:
-((4x+1)(6x-5))= 0
Multiply both sides by -1:
(4x+1)(6 x-5)= 0
Now we have:
4x+1= 0 or 6x-5= 0
Subtract 1 from both sides:
4x= -1 or 6x-5= 0
On the first equation, divide both sides by 4. On the second equation, add 5 on both sides:
x= -1/4 or 6x= 5 On the second equation, divide both sides by 6:
x= -1/4 or x= 5/6
Try these numbers and see which one makes your equation an equality. Only one will work, that will be your answer.
Re. graphing, if you don't have a graphing calculator and don't know how to graph using a spreadsheet like Excel maybe you can find a graphing calculator online.